HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ISOTOPE EFFECTS IN THE RO-VIBRATIONAL SPECTRUM OF FORMALDEHYDE

\underline{OLEKSIY A. SMOLA} , SERGEI N.YURCHENKO and JONATHAN TENNYSON, Department of Physics and Astronomy, University College London,

Gower Street, WC1E 6BT London, United Kingdom

Abstract

We present ro-vibrational line lists for three isotopologues of formaldehyde (H_{2}CO), D_{2}^{12}C^{16}O, H_{2}^{13}C^{16}O and D_{2}^{13}C^{16}O, computed using the variational nuclear-motion program TROVE (Theoretical ROVibrational Energies)^{1} in conjunction with an empirical potential energy surface (PES) by Al-Refaie et al. (2015)^{2} H2CO-AYTY. In order to take into account the effects introduced by the isotope substitutions, a mass-dependent correction to the H2CO-AYTY PES was constructed by fitting to the ro-vibrational energies of D_{2}^{12}C^{16}O, H_{2}^{13}C^{16}O and D_{2}^{13}C^{16}O, experimentally derived through the MARVEL procedure.^{3} We show that the mass-depended PES constructed using this procedure retains the quality of the parent molecule for all three isotopologues. We aim to use it for production of accurate ro-vibrational line lists for other isotopologues as part of the ExoMol project.^{4}

1 doi:10.1016/j.jms.2007.07.009S.N. Yurchenko, W. Thiel, and Per Jensen, J. Mol. Spectrosc., 245, 126 (2007)

2 doi:10.1016/j.jms.2007.07.005A. F. Al-Refaie, A. Yachmenev, J. Tennyson, and S.N. Yurchenko.

Mon. Not. R. Astron. Soc., 448, 1704 (2015)

3 doi:10.1093/mnras/stv091T. Furtenbacher, A. G. Császár and J. Tennyson, J. Mol. Spectrosc., 245, 115, (2007)

4 doi:10.1016/j.jqsrt.2020.107228J. Tennyson et al., J. Quant. Spectrosc. Radiat. Transf., 255, 107228, (2020)

p-number: p174  Submitted on Thu, 15 Jun 2023 12:15:59 +0200